The school board estimates that there will be a 2.5% increase in enrollment for next year. If the present enrollment is 8,657 what is the estimated enrollment?
Answer by stanbon(75887)

The school board estimates that there will be a 2.5% increase in enrollment for next year. If the present enrollment is 8,657 what is the estimated enrollment?
------------
Estimated enrollment = 8657 + 0.025(8657)
Ee = 8657(1+0.025)
Ee = 8657(1.025)
Ee = 8873 when rounded to a whole number.
==============
Cheers,
Stan H.